# devoted to memory of Alexei Alexandrovich Starobinsky

### v15

@@ V.6 Heat exchange between black holes in the multi-metric ensemble
VI Thermodynamics of de Sitter state and $f(R)$ gravity
VI.1 Thermodynamic variables in $f(R)$ gravity
The $T^{2}$ dependence of the de Sitter vacuum energy on temperature suggests the modification of the thermodynamic Gibbs-Duhem relation for the quantum vacuum and to the reformulation of the vacuum pressure. The conventional vacuum pressure $P_{\rm vac}$ obeys the equation of state $w=-1$ and enters the energy momentum tensor of the vacuum medium in the form:
This pressure $P_{\rm vac}$ does not satisfy the standard thermodynamic Gibbs-Duhem relation, $Ts_{\rm vac}=\epsilon_{\rm vac}+P_{\rm vac}$, because the right hand side of this equation is zero. The reason for that is that in this equation we did not take into account the gravitational degrees of freedom of quantum vacuum. Earlier it was shown, that gravity contributes with the pair of the thermodynamically conjugate variables: the gravitational coupling $K=\frac{1}{16\pi G}$ and the scalar Riemann curvature ${\cal R}$, see Refs.KlinkhamerVolovik2008c ; Volovik2022 ; Volovik2020 . The contribution of the term $K{\cal R}$ to thermodynamics is similar to the work density.Hayward1998 ; Hayward1999 ; Jacobson1995 ; Odintsov2023a
The quantities $K$ and ${\cal R}$ can be considered as the local thermodynamic variables, which are similar to temperature, pressure, chemical potential, number density, spin density, etc., in condensed matter physics. Indeed, since the de Sitter spacetime is maximally symmetric, its local structure is characterized by the scalar curvature alone, while all the other components of the Riemann curvature tensor are expressed via ${\cal R}$:
That is why the scalar Riemann curvature as the covariant quantity naturally serves as one of the thermodynamical characteristics of the macroscopic matter.Pronin1987 ; Pronin1995 Another argument is related to the so-called Larkin-Pikin effect.LarkinPikin1969 This is the jump in the number of degrees of freedom, when the fully homogeneous state is considered. One has the extra parameters, which are space independent, but participate in thermodynamics. Polyakov1991 ; KlinkhamerVolovik2008 ; Polyakov2022 The same concerns the constant electric and magnetic fields in vacuo, which add three more degrees of freedom. These constant fields are mutually independent, in contrast to the spacetime-dependent fields connected by the Maxwell equations.KlinkhamerVolovik2008 The scalar curvature ${\cal R}$ in the de Sitter vacuum, which is constant in space-time, also serves as such thermodynamic parameter. Then the gravitational coupling $K=df/d{\cal R}$ serves as the analog of the chemical potential, which is constant in the full equilibrium.
VI.2 Gibbs-Duhem relation in $f(R)$ gravity
The new thermodynamic variables, $K$ and ${\cal R}$, which come from the gravity, and Eq. (16) for the entropy density allow us to introduce the corresponding Gibbs-Duhem relation for de Sitter vacuum, which has the conventional form:
$Ts_{\rm vac}=\epsilon_{\rm vac}+P_{\rm vac}-K{\cal R}\,.$ (52)
This equation is obeyed, since $\epsilon_{\rm vac}+P_{\rm vac}=0$; ${\cal R}=-12H^{2}$; and $Ts_{\rm vac}=12\pi^{2}KT^{2}=12KH^{2}$, which supports the earlier proposal that $K$ and ${\cal R}$ can be considered as the thermodynamically conjugate variables.Volovik2022 ; Volovik2020
$P=P_{\rm vac}-K{\cal R}\,.$ (53)
Then the conventional Gibbs-Duhem relation is satisfied:
$Ts_{\rm vac}=\epsilon_{\rm vac}+P\,.$ (54)
VI.3 Entropy of cosmological horizon in terms of effective gravitational coupling
Let us show that equation $S_{\rm hor}=4\pi KA$ remains valid also in the $f({\cal R})$ gravity, but with the gravitational coupling determined as the thermodynamic conjugate to the curvature. In the $f({\cal R})$ gravity the action is:
$S=-\int d^{4}x\sqrt{-g}f({\cal R})\,.$ (55)
The generalization of the modified Gibbs-Duhem relation for the de Sitter states (i.e. for the states with constant four-dimensional curvature) in the $f({\cal R})$ gravity is:
$Ts_{\rm vac}=\epsilon_{\rm vac}+P_{\rm vac}-K{\cal R}=-K{\cal R}\,,$ (56)
$\epsilon_{\rm vac}=f({\cal R})-K{\cal R}\,\,,\,\,K=\frac{df}{d{\cal R}}\,.$ (57)
Here $K$ is the natural definition of the variable, which is thermodynamically conjugate to the curvature ${\cal R}$, while $\epsilon_{\rm vac}$ serves as the corresponding thermodynamic potential. In the equilibrium de Sitter state the curvature is determined by equation:
$2f({\cal R})={\cal R}\frac{df}{d{\cal R}}\,.$ (58)
The value of the local entropy of the de Sitter state $s_{\rm vac}$ follows from Eq.(56), assuming that the local temperature of the equilibrium dS states is $T=H/\pi$. Then the total entropy of the Hubble volume $V_{H}$ is given by the same Eq.(18):
$S_{\rm bulk}=s_{\rm vac}V_{H}=4\pi KA=S_{\rm hor}\,.$ (59)
But now $K$ is the effective gravitational coupling in Eq.(57). This generalization of the Gibbons-Hawking entropy was discussed in Refs.Odintsov2005 ; KlinkhamerVolovik2008c ; Brustein2009 ; Chao-Qiang Geng2019 . But here it was obtained using the local thermodynamics of the de Sitter vacuum. This demonstrates that the local thermodynamics of the de Sitter vacuum is valid also for the $f({\cal R})$ gravity. The effective gravitational coupling $K$ serves as one of the thermodynamic variable of the local thermodynamics. This quantity plays the role of the chemical potential, which is thermodynamically conjugate to the curvature ${\cal R}$, and it is constant in the thermodynamic equilibrium state of de Sitter spacetime.
VI.4 Example of quadratic gravity
For illustration, we consider an example of the modification of the gravitational coupling $K$ in the de Sitter environment. In the conventional Einstein gravity, where $f({\cal R})=K_{0}{\cal R}+\Lambda$, the de Sitter state has the equilibrium value of the curvature, ${\cal R}_{0}=-2\Lambda/K_{0}=-12H^{2}$. Let us add the quadratic term to the Einstein action:Odintsov2005 ; KlinkhamerVolovik2008c
$f({\cal R})=K_{0}{\cal R}-p{\cal R}^{2}+\Lambda\,.$ (60)
Then one obtains the following equations for the equilibrium value of the curvature ${\cal R}_{0}$, the entropy of the Hubble volume $S_{\rm hor}$ and the equilibrium value of the effective coupling $K$:
${\cal R}_{0}=-2\frac{\Lambda}{K_{0}}=-12H^{2}\,,$ (61)
$S_{\rm hor}=s_{\rm vac}V_{H}=4\pi KA\,,$ (62)
$K=\frac{df}{d{\cal R}}\big{|}_{{\cal R}={\cal R}_{0}}=K_{0}+p\frac{\Lambda}{K_{0}}\,.$ (63)
The equilibrium curvature in the de Sitter space ${\cal R}_{0}$ is obtained from Eq.(58). It is the same as in Einstein gravity, because the quadratic terms in Eq.(58) are cancelled. The local entropy $s_{\rm vac}$, which follows from Eq.(56), is determined by the modified gravitational coupling $K$. As a result, the entropy of the Hubble volume in Eq.(62), which we identify with the entropy of the horizon $S_{\rm hor}$, is also determined by the modified coupling $K$. The latter is given by Eq.(63).
The local entropy $s_{\rm vac}$ changes sign for $K<0$, while the cosmological expansion is still described by the de Sitter metric. However, the negative $K$ requires the negative parameter $p<0$, which marks the instability of such de Sitter vacuum.Odintsov2005
VII From de Sitter thermodynamics to de Sitter decay
VII.1 de Sitter state as thermal bath for matter
The extension of the thermodynamics to the $f({\cal R})$ gravity supports the idea that the de Sitter vacuum is the thermal state with the local temperature $T=H/\pi$. It is interesting that the gravitational temperature, which is twice the Hawking temperature, has been also obtained in the de Sitter limit in Ref.Boyle2024 , see the footnote 2 on page 4.
The nonzero local temperature of the gravitational vacuum suggests that the de Sitter vacuum is locally unstable towards the creation of thermal matter from the vacuum by thermal activation, see Section II.2. This is distinct from the mechanism of creation of the pairs of particles by Hawking radiation from the cosmological horizon, which may or may not lead to the decay of the vacuum energy. There are still controversies concerning the stability of the de Sitter vacuum caused by Hawking radiation, see e.g. Kamenshchik2022 ; Polyakov2012 ; Polyakov2022 and references therein.
VII.2 de Sitter decay due to thermalization of matter by de Sitter heat bath
To describe the decay of the vacuum due to activation and thermalization of matter, the extension of the Starobinsky analysis of the vacuum decayStarobinsky1994 ; Starobinsky1996 ; Starobinsky1997 ; Starobinsky2023 is needed. Especially the revolutionary stochastic inflation approach pioneered by StarobinskyStarobinsky1982 ; Starobinsky1986 is extremely useful, although it requires some modifications.Pattison2019 ; Cruces2022 ; Cruces2022b ; Shellard2024 This also includes the so-called separate universe approach, which is somewhat similar to the multi-metric gravity discussed in Sec. IV. Here we consider the simple phenomenological scenario based on the energy exchange between the thermal de Sitter and the thermal matter.
The thermal exchange between the de Sitter heat bath and the excited matter generates the thermal relativistic gas. The temperature of relativistic gas tends to approach the temperature $T=H/\pi$ of the de Sitter background. Then the energy density of this matter $\epsilon_{M}$ tends to approach the value $\epsilon_{M}\sim T^{4}$. In terms of the Hubble parameter, $\epsilon_{M}\rightarrow bH^{4}$, were the dimensionless parameter $b$ depends on the number of massless relativistic fields. For example, $b=7N_{F}/120\pi^{2}$ for $N_{F}$ species of massless Weyl fermions.
$\partial_{t}\epsilon_{M}=-4H(\epsilon_{M}-bH^{4})\,.$ (64)
This equation describes the tendency of matter to approach the local temperature of the vacuum, $T=H/\pi$. The extra gain of the matter energy, $4bH^{5}$, must be compensated by the corresponding loss of the vacuum energy:
$\partial_{t}\epsilon_{\rm vac}=-4bH^{5}\,,$ (65)
Here we use for simplicity the conventional general relativity with $K=K_{0}$ and $\epsilon_{\rm vac}=\Lambda=-\frac{1}{2}K{\cal R}=6KH^{2}$.
This phenomenological description of the energy exchange between the de Sitter vacuum and matter does not depend on the details of the microscopic (UV) theory, and requires only the slow-roll condition, i.e. the slow variation of the Hubble parameter, $|\dot{H}|\ll H^{2}$, or which is the same $|\dot{T}|\ll T^{2}$.
Since the vacuum energy density is $\epsilon_{\rm vac}=6KH^{2}$, one obtains from Eq.(65) the following time dependence of the Hubble parameter and of energy densities:
$H(t)=b^{-1/3}M_{\rm Pl}\left(\frac{t_{\rm Pl}}{t+t_{0}}\right)^{1/3}\,,$ (66)
Here $M_{\rm Pl}$ is the Planck mass, $M_{\rm Pl}^{2}=K$, and $t_{\rm Pl}=1/M_{\rm Pl}$ is Planck time. We assume that $t_{0}\gg t_{\rm Pl}$, and thus $\dot{H}\ll H^{2}$.
Thus the thermal character of the de Sitter state determines the process of its decay. The obtained power law decay of $H$ in Eq.(66) was also found in Refs. Padmanabhan2003 ; Padmanabhan2005 ; Markkanen2018 ; Markkanen2018a ; Roman2020 ; Gong2021 , although using different approaches. In the Padmanabhan modelPadmanabhan2003 ; Padmanabhan2005 the de Sitter horizon is considered as the photosphere with the Gibbons-Hawking temperature and with the radiative luminosity $dE/dt\propto T^{4}A_{H}$, where $A_{H}=4\pi/H^{2}$ is the area of horizon. Since the energy of the Hubble volume is $E\sim M_{\rm Pl}^{2}/H$, and $T^{4}A_{H}\sim H^{2}$, this leads to Eq.(67) for the vacuum energy density. As was mentioned by Padmanabhan,Padmanabhan2005 in his model the late time cosmological constant is independent of the initial value, see Eq.(67) at $t\gg t_{0}$.
So, the scenario of thermalization of matter by the de Sitter heat bath gives Eqs. (64) and (65). They produce the inflation in terms of two phenomenological parameters, $b$ and $t_{0}$, which determine the decay of the vacuum energy density in Eq.(67).
VII.3 Connection to holographic principle
$\frac{dV_{H}}{dt}=G\,(N_{\rm hor}-N_{\rm bulk})\,.$ (70)
Instead of the speculative degrees of freedom, we consider here the corresponding thermodynamic entropy. We already found that according to Eq.(18) the gravitational entropy of the Hubble volume is equal to the entropy related to horizon, $S_{\rm bulk}=S_{\rm hor}$. Let us consider what happens in the presence of matter. On one hand, matter adds its contribution $S_{M}$ to the bulk entropy, and thus $S_{\rm bulk}=S_{M}+S_{\rm hor}$. On the other hand, matter violates the de Sitter symmetry, which leads to the time dependence of the Hubble volume.
$s_{M}(t)=\frac{4\pi}{3}bH^{3}(t)\,.$ (71)
$S_{M}=s_{M}(t)V_{H}(t)=\left(\frac{4\pi}{3}\right)^{2}b\,.$ (72)
$\frac{dV_{H}}{dt}=\frac{4\pi}{3}bM_{\rm Pl}^{2}\,.$ (73)
Then comparing Eq.(73) with Eq.(72) and using equation $S_{M}=S_{\rm bulk}-S_{\rm hor}$ one obtains the general relation between the expansion of the Hubble volume and the difference between the bulk entropy and and its holographic surface value:
$\frac{dV_{H}}{dt}=\frac{3}{4\pi}l_{\rm Pl}^{2}S_{M}=12G\,(S_{\rm bulk}-S_{\rm hor})\,.$ (74)
Up to the sign and numerical coefficient, this coincides with Eq.(70), which supports the Padmanabhan holographic conjecture.Padmanabhan2012
VII.4 de Sitter decay and Zel’dovich stiff matter
As was mentioned by Padmanabhan,Padmanabhan2005 his model leads to a late time cosmological constant in Eq.(67), which is independent of the initial value, but its value is still far too large. Can we fix this? In Sec. VI.2 we obtained that the thermodynamics of de Sitter thermal bath corresponds to the Zel’dovich stiff matter with $w=1$. Let us try the stiff matter scenario using the phenomenological approach.
For the matter with $w=1$, the dissipative Hubble friction equation is $\partial_{t}\epsilon_{\rm stiff}=-6H\epsilon_{\rm stiff}$. In our case, the stiff matter analog is thermal, with linear in $T$ entropy density, and its temperature tends to approach the heat bath temperature $T=H/\pi$. Then instead of Eqs.(64) and (65) one obtains
$\partial_{t}\epsilon_{\rm stiff}=-6H(\epsilon_{\rm stiff}-\tilde{b}H^{2})\,,$ (75)
and
$\partial_{t}\epsilon_{\rm vac}=-6\tilde{b}H^{3}\,.$ (76)
This gives the following power law decay of the vacuum energy density:
$\epsilon_{\rm vac}(t)\sim\epsilon_{M}(t)\sim M_{\rm Pl}^{4}\left(\frac{t_{\rm Pl}}{t+t_{0}}\right)^{2}\,,$ (77)
with the reasonable values of the vacuum energy density and the energy density of matter in the present time:
$\epsilon_{\rm vac}(t_{\rm present})\sim\epsilon_{M}(t_{\rm present})\sim\frac{M_{\rm Pl}^{2}}{t_{\rm present}^{2}}\sim 10^{-120}M_{\rm Pl}^{4}\,.$ (78)
The same behaviour of the vacuum energy was obtained in Ref.KlinkhamerVolovik2008c (see also Ref. Volovik2013 ). This was also obtained in Ref.Klinkhamer2012 , where the Polyakov scenarioPolyakov2012 of the infrared instability of the de Sitter space was discussed.
[…]

### v16

@@ VII.8 Heat exchange between black holes in the multi-metric ensemble
This is the effect of thermalization.
VIII Black and white holes entropy from macroscopic quantum tunneling
VIII.1 Collective canonically conjugate variables for Schwarzschild black hole
Considering the thermodynamics of gravity systems we used the thermodynamically conjugate variables – the gravitational coupling $K=1/16\pi G$ and the scalar curvature ${\cal R}$, which as the covariant quantity may serve as one of the thermodynamical characteristics of the macroscopic matter.Pronin1987 These variables are the local thermodynamic variables, which are similar to temperature, pressure, chemical potential, number density, etc., in condensed matter physics.
The gravitational coupling $K$ is determined by the UV microscopic physics, but in the description of gravity as macroscopic phenomenon in the IR limit it is the collective variable. In the superfluid 3He-A with Weyl fermionic quasiparticles, both the coupling $K$ in the effective gravity and the fine structure “constant“ $\alpha$ in the effective electrodynamics are determined by physics on the microscopic (atomic) level.Volovik2003 In the microscopic theory, one obtains $K\propto\Delta_{0}^{2}$ and $1/\alpha\propto\ln(\Delta_{0}/T)$, where $\Delta_{0}$ is the gap amplitude and $T$ is the temperature of the liquid. The gap amplitude $\Delta_{0}$ plays the role of the ultraviolet cut-off $M_{\rm Pl}$, while $T$ provides the infrared cut-off. For the relativistic quantum fields with massless particles, the infrared cut-off is either the temperature $T$ or the strength of the fields. In the inhomogeneous superfluid (inhomogeneous vacuum) both $K$ and $\alpha$ depend on coordinates. It is not surprizing that in the relativistic quantum vacuum there is also the connection between the gravitational coupling $K$ and the coupling $\alpha$ in quantum electrodynamics as suggested in Refs. Landau1955 ; Akama ; AkamaTerazawa ; Terazawa ; TerazawaAkama ; Terazawa1981 ; KlinkhamerVolovik2005
Now, for the discussion of the quantum-mechanical tunneling of the macroscopic objects we need the collective dynamical variables, instead of the thermodynamic variables. The canonically conjugate dynamical variables, which are relevant for the black hole, are the gravitational coupling $K$ and the horizon area $A=4\pi R^{2}$. Bekenstein Bekenstein1974 proposed that $A$ is an adiabatic invariant and thus can be quantized according to the Ehrenfest principle, that classical adiabatic invariants may correspond to observables with discrete spectrum. So the area of the horizon $A$ is the proper candidate for the quantum mechanics of the black hole.
VIII.2 Modified first law of black hole thermodynamics
Let us first consider how these variables enter the black hole thermodynamics. For that it is convenient to use the redefined gravitational coupling $\tilde{K}=4\pi K=1/4G$. In terms of this coupling the Hawking temperature of Schwarzschild black hole and its Bekenstein entropy are:
$T_{\text{BH}}=\frac{\tilde{K}}{2\pi M}~{}~{},~{}~{}S_{\text{BH}}=\frac{\pi M^{2}}{\tilde{K}}=A\tilde{K}\,.$ (96)
If $\tilde{K}$ is a global thermodynamic variable, one obtains the following modification of the first law of black hole thermodynamics:
$dS_{\text{BH}}=d(A\tilde{K})=\pi d(M^{2}/\tilde{K})=-\pi\frac{M^{2}}{\tilde{K}^{2}}d\tilde{K}+2\pi\frac{M}{\tilde{K}}dM$ (97)
or
$dS_{\text{BH}}=-Ad\tilde{K}+\frac{dM}{T_{\text{BH}}}\,.$ (98)
This modification is similar to the modification in terms of the moduli fields.Gibbons1996 But in our case the thermodynamic variable, which is conjugate to the thermodynamic variable $\tilde{K}$, is the product of the black hole area and the black hole temperature, $AT_{\text{BH}}$. On the other hand in dynamics, $\tilde{K}$ and $A$ are canonically conjugate variables, see Sec.VIII.4.
In general, the variable $\tilde{K}$ is local and depends on space coordinate, but in the same way as for the moduli fields,Gibbons1996 the black hole thermodynamics is determined by the asymptotic value of $K$ at spatial infinity. In Eq.(98), $\tilde{K}\equiv\tilde{K}({\infty})$ is the global quantity, which characterizes the quantum vacuum in full equilibrium, i.e. far from the black hole.
Also, the variable $\tilde{K}$ allows us to study the transition to the vacuum without gravity, i.e. to the vacuum where $\tilde{K}\rightarrow\infty$ and thus $G\rightarrow 0$, see Sec.VIII.4.
VIII.3 Adiabatic change of coupling $K$ and adiabatic invariant
Let us change the coupling $\tilde{K}$ and the black hole mass $M$ adiabatically, i.e. at constant entropy of the black hole. Then the equation $dS_{\text{BH}}=0$ gives
$\frac{dM}{d\tilde{K}}=AT_{\text{BH}}=\frac{M}{2\tilde{K}}\,.$ (99)
This shows that $M^{2}/\tilde{K}={\rm const}$ is the adiabatic invariant for the spherical electrically neutral black hole. Thus according to the Bekenstein conjecture Bekenstein1974 , it can be quantized in quantum mechanics:
$\frac{M^{2}}{\tilde{K}}=aN\,.$ (100)
Here $N$ is integer, and $a$ is some fundamental dimensionless parameter of order unity. If this conjecture is correct, one obtains the quantization of the entropy of Schwartzschild black hole:
$S_{\text{BH}}(N)=\pi\frac{M^{2}}{\tilde{K}}=\pi aN\,.$ (101)
The Bekenstein idea on the role of adiabatic invariants in quantization of the black hole requires further consideration, see some approaches to that in Refs. Horowitz1996 ; Barvinsky2001 ; Barvinsky2002 ; Ansorg2012 ; Visser2012 ; Tharanath2013 . In particular, the similarity between the energy levels of Schwarzschild black hole and the hydrogen atom has been suggested. Bekenstein1997 ; Corda2015 We leave this problem for the future. This consideration should be supported by microscopic theory, see e.g., Ref. Carlip2014 .
In this respect, the condensed matter analogs can be useful, since in the condensed matter systems the physics is known both on macro and micro (atomic) levels. One example is provided by consideration of the quantum nucleation of the vortex ring in moving superfluids – the vortex instanton. It shows that the corresponding entropy which determines the nucleation process, $\exp(-S_{\text{ring}})$, is quantized:
$S_{\text{ring}}(N)=2\pi N\,.$ (102)
Here $N$ is the number of atoms involved in the process of the vortex instanton, see Sec. 26.4 and Eqs.(26.20)-(26.21) in Ref. Volovik2003 . This would correspond to the parameter $a=2$ in Eq.(100).
VIII.4 $A$ and $K$ as canonically conjugate variables and black-hole – white-hole quantum tunneling
The canonically conjugate variables $A$ and $K$ allow us to consider the quantum mechanical tunneling from the black hole to the white hole – the process discussed in Refs. Barcelo2014 ; Barcelo2017 ; Rovelli2018 ; Rovelli2019 ; Rovelli2018b ; Uzan2020b ; Uzan2020c ; Uzan2020d ; Bodendorfer2019 ; Rovelli2021a and in references therein. In the semiclassical description of this macroscopic quantum tunneling, the trajectory in the $(\tilde{K},A)$ phase space must be considered, which in the complex plane connects the black and white holes.
Note the difference from consideration in Sec. VIII.3, where $\tilde{K}$ varies in the adiabatic regime, i.e. at fixed entropy $S_{\rm BH}$, while the area $A$, temperature $T_{\rm BH}$ and mass $M$ follow the variation of $\tilde{K}$. In the dynamic regime, which is relevant for the description of quantum tunnelling, the parameter $\tilde{K}$ varies at fixed energy (fixed mass $M$ of the black hole), while the area $A$, the temperature $T_{\rm BH}$ and entropy $S_{\rm BH}$ follow the variation of $\tilde{K}$.
As in the case of the semiclassical consideration of the Hawking radiation in terms of the quantum tunneling, we shall use the Painleve-Gullstrand coordinate system with the metric:
$ds^{2}=-dt^{2}(1-{\bf v}^{2})-2dt\,d{\bf r}\cdot{\bf v}+d{\bf r}^{2}\,,$ (103)
where for the Schwartzschild black hole one has
${\bf v}({\bf r})=\mp\hat{\bf r}\sqrt{\frac{R}{r}}=\mp\hat{\bf r}\sqrt{\frac{M}{2r\tilde{K}}}=\mp\hat{\bf r}\sqrt{\frac{2MG}{r}}\,,$ (104)
where $R$ is the radius of the horizon; the minus sign corresponds to the black hole and the plus sign describes the white hole. In the theory with the variable gravitational coupling, the sign changes at the singularity $\tilde{K}=\infty$ (i.e. at $G=0$), when the black hole shrinks to a point and then expands as a white hole. The point $\tilde{K}=\infty$, where gravity disappears, serves as the branch point, where the velocity of the freely falling observer changes sign.
The vector ${\bf v}$ is normal to the surface of the horizon. When ${\bf v}$ changes sign, the horizon area, $A$, also changes sign: it crosses zero at $\tilde{K}=\infty$ and becomes negative on the white-hole side of the process, $A\rightarrow-A$. This also could mean that due to connection between the area and entropy the white hole may have negative entropy. This is what we shall discuss in the section VIII.5.
The quantum tunneling exponent is usually determined by the imaginary part of the action on the trajectory, which transforms the black hole to white hole. In terms of Euclidean action one has:
[…]
Here the semiclassical trajectory $C$ is at $M=\rm{const}$, and thus $A(\tilde{K}^{\prime})=\pm\pi M^{2}/(\tilde{K}^{\prime})^{2}$. Along this trajectory the variable $\tilde{K}^{\prime}$ changes from $\tilde{K}$ to the branch point at $\tilde{K}^{\prime}=\infty$, and then from $\tilde{K}^{\prime}=\infty$ to $\tilde{K}^{\prime}=\tilde{K}$ along the other branch, where the area $A(\tilde{K}^{\prime})<0$. The integral gives the tunneling exponent of the transition to the white hole
$I_{\text{BH}\rightarrow\text{WH}}=\,2\pi M^{2}\int_{\tilde{K}}^{\infty}\frac{d\tilde{K}^{\prime}}{\tilde{K}^{\prime 2}}=2\pi\frac{M^{2}}{\tilde{K}}\,.$ (106)
The tunneling exponent in Eq.(106) can be expressed in terms of the black hole entropy in Eq (96), and for the probability of transition one obtains:
$p_{\text{BH}\rightarrow\text{WH}}\propto\exp{\left(-2\pi M^{2}/\tilde{K}\right)}=\exp{\left(-2S_{\text{BH}}\right)}\,.$ (107)
It is important that Eq.(107) contains exactly twice the black hole entropy. This allows us to consider again the entropy of the white hole.
VIII.5 Negative entropy of white hole
The connection between the probability of the transition and the thermodynamic fluctuationsLandau_Lifshitz is applicable also to the transition between the black and white holes. The total change of the entropy in this process is $\Delta S=S_{\text{WH}}-S_{\text{BH}}$. According to Eq.(107) this change is equal to $-2S_{\text{BH}}$. Then, from equation $S_{\text{WH}}-S_{\text{BH}}=-2S_{\text{BH}}$, one obtains that the entropy of the white hole is equal with the opposite sign to the entropy of the black hole with the same mass:
$S_{\text{WH}}(M)=-S_{\text{BH}}(M)\,.$ (108)
This means that the white hole, which is obtained by the quantum tunneling from the black hole and thus has the same mass $M$ as the black hole, has the negative temperature $T_{\text{WH}}=-T_{\text{BH}}$ and the negative area $A_{\text{WH}}=-A_{\text{BH}}$. According to the first law in Eq. (98) applied now to the white hole, this gives the negative entropy, $S_{\text{WH}}=-S_{\text{BH}}$.
Such anti-symmetry between the black and white holes is similar to the anti-symmetry between the expanding and contracting de Sitter state discussed in Sec. VII.2.
The discussed transition from the black hole to the white hole with the same mass $M$ is not the thermodynamic transition. It is the quantum process of tunneling between the two quantum states. It is one of many routes of the black hole evaporation, including formation of small white hole on the late stage of the decay. Barcelo2014 ; Barcelo2017 ; Rovelli2018 ; Rovelli2019 ; Rovelli2018b ; Uzan2020b ; Uzan2020c ; Uzan2020d ; Bodendorfer2019 ; Rovelli2021a The uniqueness of this particular route, the hidden information and the (anti)symmetry between the black and white holes are combined to produce the negative entropy of the white hole.
VIII.6 Black hole to white hole transition as series of the Hawking radiation co-tunneling
Let us show that the result (107), where the probability of the quantum tunneling between the black and white holes transition is determined by the twice the black hole entropy, is supported by consideration of the conventional Hawking radiation of particles from the black hole. The tunneling exponent of radiation can be expressed in terms of the change in the entropy of the black hole after radiation, $p\propto e^{\Delta S_{\text{BH}}}$, see Refs. Kraus1997 ; Parentani1997 ; Berezin1999 ; Wilczek2000 . This demonstrates that the quantum process of Hawking radiation can be considered as thermodynamic fluctuation.Landau_Lifshitz
We consider the process of the co-tunneling, in which the particle escapes the black hole by quantum tunneling, and then this particle tunnels to the white hole through the white hole horizon (this is the analog of the electron tunneling via an intermediate virtual state in electronic systems Feigelman2005 ; Glazman2005 ). This process takes place at the fixed total mass $M$. The tunneling exponent for this process to occur is $e^{(\Delta S_{\text{BH}}+\Delta S_{\text{WH}})}$. Summation of all the processes of the tunneling of matter from the black hole to the formed white hole gives finally Eq.(107):
$p\propto e^{\sum(\Delta S_{\text{BH}}+\Delta S_{\text{WH}})}=e^{2\sum\Delta S_{\text{BH}}}=\exp{\left(-2S_{\text{BH}}\right)}\,.$ (109)
Here we took into account the (anti)symmetry in the dynamics of black and white holes in the process of quantum tunneling, $\sum\Delta S_{\text{BH}}=\sum\Delta S_{\text{WH}}$.
VIII.7 Emission of small black holes vs Hawking radiation
The principle that the macroscopic quantum tunneling can be considered as thermodynamic fluctuation can be also applied to the process of the creation of pairs black holes,HawkingHorovitz1995 to the process of splitting of the black hole into two or several smaller black holes with the same total mass, see e.g. Ref.HyeyounChung2011 , and to other processes with macroscopic objects. For example, the probability of the decay of the black hole with mass $M$ into two black holes with $M_{1}+M_{2}=M$ is:
$p(M\rightarrow M_{1}+M_{2})\propto\,$ (110)
$e^{S_{\text{BH}}(M_{1})+S_{\text{BH}}(M_{2})-S_{\text{BH}}(M_{1}+M_{2})}=e^{-8\pi GM_{1}M_{2}}.$ (111)
In the particular limit case $m=M_{1}\ll M_{2}\approx M$, this channel of the black hole decay describes the emission of the small black hole with mass $m$ by the large black hole with mass $M$:
$p(m,M)\propto\exp{\left(-\frac{m}{T_{\rm BH}(M)}\right)}\,\,,\,\,m\ll M\,.$ (112)
This shows that the macroscopic tunneling process of emission of a small black hole by a large black hole is governed by the same Hawking temperature $T_{\rm BH}(M)=1/(8\pi GM)$ as the Hawking radiation of a particle, which tunnels across the horizon.
However, there is the difference. In Eq.(112) the quadratic term $m^{2}$ is neglected. In general case, one obtains from Eq.(111):
$p(m,M-m)\propto\exp{\left(-8\pi Gm(M-m)\right)}\,.$ (113)
This equation demonstrates the effect of back reaction – the correction to the Hawking radiation caused by the reduction of the black hole mass after the radiation of a small black hole. The similar correction due to the back reaction in the process of radiation of particles was obtained by Parikh and Wilczek Wilczek2000 (see also Ref.Kraus1997 ):
$p(\omega,M-\omega)\propto\exp{\left(-8\pi G\omega\left(M-\frac{\omega}{2}\right)\right)}\,,$ (114)
where $\omega$ is the energy of the emitted particle. As distinct from Eq.(113) for emission of a small black hole, the Eq.(114) contains the factor $1/2$. The reason for such difference is that in case of the emission of a small black hole, the probability of emission contains the extra term compared to the emission of particles – the entropy of the emitted black hole, $S(m)=4\pi Gm^{2}$:
$p(m,M-m)\propto\exp{\left(-8\pi Gm\left(M-\frac{m}{2}\right)+4\pi Gm^{2}\right)}\,.$ (115)
As a result the Eq.(113) is restored. This also coincides with the result for the radiation of the self-gravitating shell.KrausWilczek1995
Eq.(111) can be extended to emission of several black holes. For example, the probability of emission of $N$ black holes with masses $m=M/N$ is:
$p(M\rightarrow mN)\propto\exp{\left(-4\pi GM^{2}\left(1-\frac{1}{N}\right)\right)}\,.$ (116)
For large $N$ this corresponds to the probability of the destruction of the black hole by explosion within the particular channel:
$p_{N\rightarrow\infty}\propto\exp(-S_{\rm BH})\,.$ (117)
There are many channels of the destruction of the black hole in quantum tunnelling. The black hole entropy $S_{\rm BH}$ can be estimated using the probability of the black hole explosion in a single quantum event. Then, the Eq.(117) suggests that the entropy of the black hole is determined by the number of the possible channels leading to the destruction of the black hole.
The discussed scenario of the destruction of black holes by explosion to the small Planck scale objects also explains the origin of the negative entropy of the white hole. The explosion is a rare process in which the entropy is reduced to zero. However, after the explosion occurs, the even the more rare process may follow, when the small Planck scale objects are collected back forming the white hole. The rarity of this process reduces the entropy even further, which leads to the negative entropy of the formed white hole. This is the reason why the white hole has the negative entropy. The latter is also supported by the consideration of super-rare process of the quantum tunnelling from the black to the white hole of the same mass in Sec. VIII.5. The decrease of the entropy in this super-rare process corresponds to the loss of the double entropy of the black hole, which results in the negative entropy of the white hole.
The negative temperature of the white hole is also not very surprising. In general, the negative absolute temperatures are consistent with equilibrium thermodynamics. All the thermodynamic properties, such as thermometry, thermodynamics of cyclic transformations, ensemble equivalence, fluctuation-dissipation relations, response theory, the transport processes and symmetry breaking phase transitions, can be reformulated to include the negative temperatures.Baldovin2021 ; Lounasmaa1997 ; Volovik2021
IX Conclusion
The starting point of our consideration was that matter immersed in the de Sitter vacuum percieves this vacuum as the heat bath with the local temperature $T=H/\pi$, where $H$ is the Hubble parameter. This temperature has no relation to the cosmological horizon, and to the Hawking radiation from the cosmological horizon. However, it is exactly twice the Gibbons-Hawking temperature, $T_{\rm GH}=H/2\pi$. The reason for such relation is the specific symmetry of the de Sitter space-time, which is similar to the invariance under translations in the Minkowski vacuum.
[…]